Crane Lumber, Inc., is considering purchasing a new wood saw that costs $50,000. The saw will generate revenues of $100,000 per year for five years. The cost of materials and labor needed to generate these revenues will total $60,000 per year, and other cash expenses will be $10,000 per year. The machine is expected to sell for $3,100 at the end of its five-year life and will be depreciated on a straight-line basis over five years to zero. Crane’s tax rate is 34 percent, and its opportunity cost of capital is 17.50 percent.

What is the project's NPV? (Do not round intermediate calculations. Round final answer to the nearest whole dollar, e.g. 5,275.)
Answer #1

rate positively ..

Computation of NPV
year 0 1 2 3 4 5
A Initial investment -50000
Operating cash flow
i sales            100,000                          100,000              100,000           100,000           100,000
ii Cost of material              60,000                            60,000                60,000             60,000             60,000
iii Other cost              10,000                            10,000                10,000             10,000             10,000
iv depreciation         10,000.00                       10,000.00           10,000.00        10,000.00        10,000.00
v=i-ii-iii-iv Profit before tax         20,000.00                       20,000.00           20,000.00        20,000.00        20,000.00
vi=v*34% Tax @ 34%           6,800.00                         6,800.00             6,800.00          6,800.00          6,800.00
vii=v-vi Net income         13,200.00                       13,200.00           13,200.00        13,200.00        13,200.00
B=vii+iv Operating cash flow=         23,200.00                       23,200.00           23,200.00        23,200.00        23,200.00
C Salvage value 2046
3100*(1-34%)
D=A+B+C Net cash flow -$50,000.00 $23,200.00 $23,200.00 $23,200.00 $23,200.00 $25,246.00
E PVIF @ 17.5% 1 0.85106383 0.724309642 0.616433738 0.524624458 0.4464889
F=D*E Present value -$50,000.00 $19,744.68 $16,803.98 $14,301.26 $12,171.29 $11,272.06 $24,293.27
NPV = $24,293.27
ans =     24,293.2735